OpenStudy (anonymous):

Given the range {-24, 6, 26}, what is the domain for the relation y = 5x - 4?

OpenStudy (aravindg):

range means the y values we get

OpenStudy (aravindg):

y=5x-4 -24=5x-4 6=5x-4 26=5x-4

OpenStudy (aravindg):

Solve for x in the 3 equations above you get the domain.